# Maneuvering measurement-coherence into measurement-entanglement

### v2

@@ Abstract
Are there relations between the two static resource powers of a quantum channel? In general, there is no relation between a channel’s state-resource power and its measurement-resource power. Consider a preparation channel $\mathcal{E}_{A}(\rho_{A})=\operatorname{Tr}_{A}(\rho_{A})\left|+\right\rangle\!\left\langle+\right|_{A}$. This quantum channel has the maximum state-cohering power as $C_{g}(\mathcal{E}_{A})=1$ but no measurement-cohering power as $C(\mathcal{E}_{A})=0$. Another quantum channel $\mathcal{G}_{A}(\rho_{A})=\langle+\rvert\rho_{A}\lvert+\rangle_{A}\left|0\right\rangle\!\left\langle 0\right|_{A}+\langle-\rvert\rho_{A}\lvert-\rangle_{A}\left|1\right\rangle\!\left\langle 1\right|_{A}$ shows the opposite case where it can generate no state-coherence as $C_{g}(\mathcal{G}_{A})=0$ while it can generate the maximum measurement-coherence $C(\mathcal{G}_{A})=1$. However, for a unital quantum channel, we establish a connection between state-cohering power and measurement-cohering power:
@@
For entanglement, we find a similar relation. The state-entangling power of a quantum channel $\mathcal{E}_{AB}$  [^2] The state-entangling power has also been used to characterize the average state-entanglement generated from arbitrary product states as in [36]. We follow the convention of [41] for consistency with the state-cohering power. is defined as
$E_{g}(\mathcal{E}_{AB})\coloneqq\max_{\sigma\in\mathbf{SepD}(A\widetilde{A}:B\widetilde{B})}E_{R}(\mathcal{E}_{AB}(\sigma_{A\widetilde{A}B\widetilde{B}})),$ (18)
@@
Conclusion.—We have established resource theories quantifying the measurement-cohering and measurement-entangling powers of quantum channels, demonstrating a conversion between these two measurement-related dynamical resources. This result completes the conceptual framework linking quantum coherence and quantum entanglement at both static and dynamic levels, as illustrated in Table 1. Specifically, while previous studies addressed the connection between quantum channel’s state-resource powers, our work connects the measurement-cohering power of a quantum channel to the measurement-entangling power. Our results integrate and extend prior studies on the measurement resources by treating quantum measurements as channels producing classical data. This approach reproduces known results on conversion from measurement-coherence to measurement-entanglement, placing them within our broader theoretical framework. In achieving our main results, we utilized the newly characterized structure of an incoherent measurement as an incoherent basis measurement followed by a classical post-processing channel.
State-resource Measurement-resource
Static [Uncaptioned image] [Uncaptioned image]
Dynamic [Uncaptioned image] [Uncaptioned image]
Table 1: Illustration of the optimal conversion of quantum coherence into entanglement at different levels of resource theory. The table categorizes these conversions into static (quantum states and measurements) and dynamic (quantum channels) scenarios. The diagrams, arranged from the top left to the bottom left in a clockwise direction, depict: the optimal conversion of quantum coherence of a quantum state $\rho_{A}$ to entanglement of a quantum state $\widetilde{\rho}_{AB}$ [19], the optimal conversion of measurement-coherence of a quantum measurement $\mathcal{M}_{A}$ to measurement-entanglement of a quantum measurement ~$\widetilde{\mathcal{M}}_{AB}$ [35], the optimal conversion of measurement-cohering power of a quantum channel $\mathcal{E}_{A}$ to measurement-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ (this work), and the optimal conversion of state-cohering power of a quantum channel $\mathcal{E}_{A}$ to state-entangling power of a quantum channel ~$\widetilde{\mathcal{E}}_{AB}$ [43].
Moreover, we demonstrated that the state-cohering power of a unital quantum channel is equivalent to the measurement-cohering power of its adjoint channel, and that a similar equivalence holds between the state-entangling power of a unital bipartite quantum channel and the measurement-entangling power of its adjoint channel.
Despite significant advances, fundamental open questions persist. While coherence-to-entanglement conversion is well-documented, the reverse transformation remains poorly understood due to the basis-independent nature of entanglement, making such conversions inherently nontrivial. Moreover, the interplay between quantum coherence and multipartite entanglement remains largely unexplored [21], partly owing to the incomplete theoretical characterization of multipartite entanglement itself.
Another critical point is that static resource generation alone cannot capture the full complexity of quantum dynamics concerning resource utilization. Indeed, quantum dynamics itself can be viewed as a resource under appropriate restrictions on quantum superchannels [53, 54, 55, 56]. For example, the generalized Hadamard gate has been investigated as a dynamic coherence resource [57], while the SWAP gate functions as a dynamic entanglement resource capable of simulating arbitrary bipartite quantum channels [42, 56, 58], analogous to how Bell states serve as universal resources for state preparation. This viewpoint connects naturally to quantum simulation frameworks [59, 60, 61], one of quantum computing’s most promising applications, emphasizing the essential relationship between experimental capabilities and required quantum resources. However, the connections among these dynamic quantum resources remain less understood compared to their static counterparts. A deeper insight into these dynamical relationships holds potential for developing more efficient algorithms, particularly in quantum simulation.
Addressing these open issues will significantly enhance our fundamental understanding of quantum resources and promote the advancement of resource-efficient quantum information processing technologies.
Acknowledgements.

### v3

@@ 5 State-resource power versus measurement-resource power
Are there relations between the two static resource powers of a quantum channel? In general, there is no relation between a channel’s state-resource power and its measurement-resource power. Consider a qubit preparation channel $\mathcal{E}_{A}(\rho_{A})=\operatorname{Tr}_{A}(\rho_{A})\left|+\right\rangle\!\left\langle+\right|_{A}$ with $\lvert+\rangle_{A}=(\lvert 0\rangle_{A}+\lvert 1\rangle_{A})/\sqrt{2}$. While $\mathcal{E}_{A}$ has the maximum state-cohering power ($C_{g}(\mathcal{E}_{A})=1$), it has zero measurement-cohering power ($C(\mathcal{E}_{A})=0$) because the effective measurement $\mathcal{I}_{A}\circ\mathcal{E}_{A}$ yields a trivial measurement whose POVM element is $I/2$. Another quantum channel $\mathcal{G}_{A}(\rho_{A})=\langle+\rvert\rho_{A}\lvert+\rangle_{A}\left|0\right\rangle\!\left\langle 0\right|_{A}+\langle-\rvert\rho_{A}\lvert-\rangle_{A}\left|1\right\rangle\!\left\langle 1\right|_{A}$ shows the opposite case, where it can generate no state-coherence as $C_{g}(\mathcal{G}_{A})=0$ while it can generate the maximum measurement-coherence $C(\mathcal{G}_{A})=1$. Meanwhile, the Hadamard gate possesses the maximum state-cohering power (converting incoherent basis states into maximally coherent states) and also the maximum measurement-cohering power (converting the incoherent measurement $\mathcal{I}$ into a measurement in the Fourier basis). However, for unital quantum channels, we establish a connection between the state-cohering power and the measurement-cohering power:
@@
For entanglement, we find a similar relation. The state-entangling power of a quantum channel $\mathcal{E}_{AB}$ [^2] The term state-entangling power has also been used to characterize the average state-entanglement generated from arbitrary product states as in Ref. [38]. We follow the convention of Ref. [44] for consistency with the state-cohering power. is defined as
$),E_{g}(\mathcal{E}_{AB})\coloneqq\max_{\sigma\in\mathbf{SepD}(A\widetilde{A}:B\widetilde{B})}E_{R}(\mathcal{E}_{AB}(\sigma_{A\widetilde{A}B\widetilde{B}}$)), (17)
@@
6 Conclusion
We have developed resource theories for quantifying measurement-cohering and measurement-entangling powers of quantum channels, thereby extending the landscape of dynamical quantum resources to include measurement-based phenomena. Our results establish a dynamical correspondence between coherence and entanglement on the measurement side: the measurement-cohering power of any quantum channel can be fully converted into its measurement-entangling power using only free operations and a dephasing step. This conversion theorem completes the coherence–entanglement correspondence at the dynamical level and mirrors the known relationship between state-cohering and state-entangling powers, as illustrated in Table 1.
A useful ingredient enabling these results is our structural characterization of incoherent measurements, showing that every incoherent POVM reduces to the incoherent basis measurement followed by classical post-processing. This structural insight allows measurement channels to be treated analogously to state preparations, permitting a unified treatment of their dynamical generation under quantum channels. Our framework also reproduces the static results of measurement-coherence and measurement-entanglement when a measurement is regarded as a channel with classical outputs.
Furthermore, we have shown that for unital channels, dynamical resource powers for states and measurements are tightly linked: the state-cohering power of a channel equals the measurement-cohering power of its adjoint map, and an analogous equivalence holds for state-entangling versus measurement-entangling powers. These dualities highlight a fundamental symmetry between state- and measurement-based quantum resources, suggesting that both arise from a single underlying structure governed by the dynamical action of quantum channels.
Despite these advances, several fundamental questions remain open. Although coherence-to-entanglement conversion is now understood in both static and dynamical regimes, it remains unclear whether all forms of entanglement can be obtained from coherence. Determining when, and to what extent, entanglement can be generated from coherence in a resource-theoretically consistent manner is an intriguing challenge. Moreover, the relationship between coherence and multipartite entanglement is still insufficiently explored [19], reflecting broader gaps in our understanding of multipartite entanglement itself.
Another important point is that static resource generation alone cannot capture the full complexity of quantum dynamics concerning resource utilization. Indeed, quantum dynamics itself can be viewed as a resource under appropriate restrictions on quantum superchannels [60, 61, 62, 63, 64]. Recent works have shown that gates such as generalized Hadamard or SWAP can serve as fundamental dynamical resource units [46, 48, 45, 63, 64], analogous to maximally coherent states or maximally entangled states in static resource theories. However, the connections among these dynamical quantum resources remain less understood compared to their static counterparts. Clarifying the relationship between dynamical coherence and dynamical entanglement resources may also prove valuable for quantum technologies, where the ability to manipulate channel-level resources underlies tasks such as fault-tolerant gate synthesis and quantum simulation. A unified understanding of these dynamical resources could help identify operations that are optimal or universal for emerging quantum devices.
Overall, our results place measurement-based dynamical resources on equal footing with their state-based counterparts and provide a unified theoretical foundation linking coherence and entanglement across static and dynamical levels. We expect that this framework will inform future advances in the broader development of resource-efficient quantum technologies, such as quantum learning and quantum simulation.